OpenStudy (anonymous):

if u=k(3,-1,0), where u is a unit vector and k>0, what is the value of k?

OpenStudy (anonymous):

\[k = \frac{ 1 }{ \sqrt{10} }\]Because we know that a unit vector has length 1. The length of \[(3,-1,0)\]is\[\sqrt{3^2 + (-1)^2 + 0^2}=\sqrt{10}\]we need to divide our vector by that to get it to be of length 1.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes, thanks so much i was stuck on the question for a while lol